OpenStudy (anonymous):

the density of smoke from a diesel truck is related to its engine speed, and may be approximated by the formula D=2r^2-14r+30. In the formula , D is measured in millions of particles per cubic foot, and are is in revolutions per minute. If the density of the smoke from a truck is 14 , what is its engine speed?

OpenStudy (anonymous):

you need to plug in 14 for D in the given equation and solve for r

OpenStudy (anonymous):

2 r^ 2 -14r +30 = 14

OpenStudy (anonymous):

2 r ^2 - 14r +16 = 0, that is we subtract 14 from both sides of the equation

OpenStudy (anonymous):

now 2 is a common factor , therefore we get, r ^2 - 7r +8 = 0

OpenStudy (anonymous):

we need to find two numbers whose sum ids -7 and product is 8

OpenStudy (anonymous):

ok but that would leave 8 revolutions per minute when it should st least be in the hundreds